OpenStudy (anonymous):

Consider g(x)=3sin2x ,write down the period of g

OpenStudy (anonymous):

period of \(\sin(bx)\) is always \(\frac{2\pi}{b}\)

OpenStudy (anonymous):

in your case you get \(\frac{2\pi}{2}=\pi\) as the period

OpenStudy (anonymous):

Is the 3 irrelevant when it comes to finding the period?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that gives you the "amplitude"

OpenStudy (anonymous):

since sine goes from -1 to 1, \(3\sin(x)\) goes from -3 to 3

OpenStudy (anonymous):

allright, thank you for your help!!! much appreciated

OpenStudy (anonymous):

sine does everything between 0 and \(2\pi\) for the input if \(2x=0\) then \(x=0\) and if \(2x=2\pi\) then \(x=\pi\) so this goes through one period from \(0\) to \(\pi\)

OpenStudy (anonymous):

yw